OpenStudy (anonymous):

Graph y=2/(x+5)-3 and the parent function. Identify horizontal asymptotes. Identify vertical asymptotes. Describe the transformations.

OpenStudy (anonymous):

i am plotting the equation using word documents,i will attach the file soon

OpenStudy (anonymous):

here is the graph

OpenStudy (anonymous):

as you can see the function is shifted by 2 units in the negative x-axis, because x+5-3=2

OpenStudy (anonymous):

you may compare this with f(x)=1/x , if you want i can draw it for you

OpenStudy (anonymous):

How do I find the asymptotes though?

OpenStudy (anonymous):

as you can see the vertical line passing through (2,0)

OpenStudy (anonymous):

can you see the vertical blue line

OpenStudy (anonymous):

look at the curve is approaching the vertical line, it should never touch that line

OpenStudy (anonymous):

this is how this specific function behaves by nature

OpenStudy (anonymous):

so you need to go by its rules

OpenStudy (anonymous):

have a look at f(x)=1/x

OpenStudy (anonymous):

you can see even the y axis has been multiplied by 2

OpenStudy (anonymous):

So it was only a shift?

OpenStudy (anonymous):